Chapter 15, Problem 064 GO The inside of a Carnot refrigerator is maintained at a temperature of 277 K, while the temperaturehelp please
Answer #1

Here,

Tc = 277 K

Th = 293 K

W = 3757 K

Qc = ?

for the Carnot cycle

W/Qc = (Th - Tc)/Tc

3757/Qc = (293 - 277)/277

solving for Qc

Qc = 65043 J

the heat that can be taken out is 65043 J